# Probability calibration for precipitation nowcasting

### v1

@@ 1 Introduction
In this note, we address the miscalibration of probabilistic NWMs in precipitation nowcasting. We introduce the expected thresholded calibration error (ETCE) as a more appropriate metric for assessing miscalibration in our setting. We study post-processing techniques designed to adjust model confidence values to better match observed precipitation frequencies. While there is active research on the topic, existing calibration methods are limited—particularly in computer vision—and, to our knowledge, absent in the context of NWMs 9710154 ; wang10205155 ; ding_local .
@@ 2.1 Expected thresholded calibration error
where $w_{b}$ are bin weights. In this application, we apply uniform weighting $w_{b}=1/B$ as precipitation is a rare event and weighting by the number of samples would emphasize dry events in the metric. In our experiments we have enough data to select $B=20$ (See Figure A.1 in the Appendix).
@@ Appendix A Appendix
A.1 Number of predictions in different confidence bins
@@
A.2 Selective scaling with Segformer classifiers
@@
A.3 Calibration effect on ETCE

### v2

@@ 1 Introduction
In this note, we address the miscalibration of probabilistic NWMs in precipitation nowcasting. We introduce the expected thresholded calibration error (ETCE) as a more appropriate metric for assessing miscalibration in our setting. We study post-processing techniques designed to adjust model confidence values to better match observed precipitation frequencies. While there is active research on the topic, existing calibration methods are limited—particularly in computer vision—and, to our knowledge, absent in the context of NWMs wang10205155 ; 9710154 ; ding_local .
@@ 2.1 Expected thresholded calibration error
where $w_{b}$ are bin weights. We expand the confidence and accuracy terms in the appendix (see Section A.1). In this application, we apply uniform weighting $w_{b}=1/B$ as precipitation is a rare event and weighting by the number of samples would emphasize dry events in the metric. In our experiments we have enough data to select $B=20$ (See Figure A.1 in the Appendix).
@@ Appendix A Appendix
A.1 ETCE details
Here, we give the full details on the confidence and accuracy terms of Eq. 2. In our dataset we have input-target pairs $(x_{i},r_{i})$ where $x_{i}$ is the DNN input and $r_{i}$ is the ground truth precipitation rate of sample $i$. Then, the term $\operatorname{conf}(b,R_{k})$ for confidence bin $b$ and precipitation threshold $R_{k}$ is given by
$\operatorname{conf}(b,R_{k})\coloneq\frac{1}{\left|B_{b}\right|}\sum_{i\in B_{b}}s_{x_{i}}(R_{k}),\quad\text{where}\ B_{b}\coloneq\left\{i:s_{x_{i}}(R_{k})\in b\right\},$ (5)
where $s_{x_{i}}(\cdot)$ is the survivability function obtained from the predicted probability for input $x_{i}$. That is, $s_{x_{i}}(R_{k})=1-\operatorname{CDF}_{\operatorname{DNN}(x_{i})}(R_{k})$.
The accuracy term is
$\operatorname{acc}(b,R_{k})\coloneq\frac{1}{\left|B_{b}\right|}\sum_{i\in B_{b}}\mathbbm{1}(r_{i}\geq R_{k}),$ (6)
where $r_{i}$ is the ground truth precipitation rate for sample $i$.
A.2 Number of predictions in different confidence bins
@@
A.3 Selective scaling with Segformer classifiers
@@
A.4 Calibration effect on ETCE